OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4.

OpenStudy (anonymous):

If a line is parallel to the equation y=-x-4, then that means it will have the same slope as that equation, but with a different y-intercept. Meaning: A line that is parallel to the equation y=2x would be y=2x+1 or y=2x+100 or y=2x-1. Just as long as it has the same coefficient before x in the original equation, you'll be good to go. In this question, you have: y=-x-4 which is the same as y=(-1)x-4 the coefficient is (-1), so just make a new equation with the same slope (same coefficient), but instead of subtracting 4 afterwards, there'll be a different number subtracted, or even added.